OpenStudy (fanduekisses):

How do I find the point slope when I only have a slope and a y-intercept?

OpenStudy (fanduekisses):

for ex: slope= -5, y-intercept=-8

OpenStudy (anonymous):

\(y\) intercept \(-8\) means the point is \((0,-8)\)

OpenStudy (anonymous):

you don't need the "point - slope" formula if you know the slope is \(m\) and the \(y\) intercept is \(b\) the equation for the line is \(y=mx+b\)

OpenStudy (fanduekisses):

OHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHHH

OpenStudy (fanduekisses):

So the x intercept is 0

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the \(x\) intercept is where the line crosses the \(x\) axis

OpenStudy (anonymous):

but if i know the slope is \(2\) and the \(y\) intercept is \(5\) then i know the equation is \[y=2x+5\]

OpenStudy (fanduekisses):

I mean if I wear to set up the point slope, it's be: (y-(-8)=-5(x-0) ?

OpenStudy (fanduekisses):

it'd*

OpenStudy (anonymous):

you could, but you wouldn't

OpenStudy (anonymous):

you would go right to \(y=-5x-8\)

OpenStudy (fanduekisses):

Oh ok, but they are asking me for a point slope

OpenStudy (anonymous):

k, then you would write what you write, or \(y+8=-5x\)

OpenStudy (fanduekisses):

oh ok

OpenStudy (fanduekisses):

Thank you :D